trentin:

Solve for x: −3|x + 7| = −12

dude:

Divide by -3 on both sides \[|x+7|=4\]

dude:

So same idea as an earlier question \[x+7=-4\quad \quad \mathrm{or}\quad \:\quad \:x+7=4\]

dude:

#1 \[x+7=−4\] Subtract 7 on both sides \[\x=-11]

dude:

#2 \[x+7=4\] Subtract 7 on both sides \[x=-3\]

trentin:

-x-7=4

trentin:

-x=4=7

trentin:

4+7 = 11

trentin:

x - 11

dude:

I don't understand where you got that -x from

trentin:

Break down the problem into these 2 equations

trentin:

Solve the 1st equation: x+7=4

trentin:

Subtract 7 from both sides

dude:

Yes so \[x=-11\]

trentin:

yes

trentin:

x=−11,−3

dude:

Yes

trentin:

can you help me on the next one